# envname-P envname#1

### v2

@@ Abstract
The fidelity with which neural networks can now generate content such as music presents a scientific opportunity: these systems appear to have learned implicit theories of such content’s structure through statistical learning alone. This offers a potentially new lens on theories of human-generated media. When internal representations align with traditional constructs (e.g. chord progressions in music), they show how such categories can emerge from statistical regularities; when they diverge, they expose limits of existing frameworks and patterns we may have overlooked but that nonetheless carry explanatory power. In this paper, focusing on music generators, we introduce a method for discovering interpretable concepts using sparse autoencoders (SAEs), extracting interpretable features from the residual stream of a transformer model. We make this approach scalable and evaluable using automated labeling and validation pipelines. Our results reveal both familiar musical concepts and coherent but uncodified patterns lacking clear counterparts in theory or language. As an extension, we show such concepts can be used to steer model generations. Beyond improving model transparency, our work provides an empirical tool for uncovering organizing principles that have eluded traditional methods of analysis and synthesis. [^1] https://musicdiscovery.media.mit.edu/
Figure 1: Multi-stage pipeline for discovering and steering interpretable concepts in generative music models. (1) Music from a large corpus is passed through a pre-trained generator to extract activations from multiple layers. (2) Sparse autoencoders reconstruct activations (also usable for steering), and features are filtered to retain the most viable candidates. (3) Retained features are characterized via musical examples and labeled using generative labeling with a multimodal LM and classifier-based labeling with pre-trained models.
@@ 1 Introduction
    A general pipeline for unsupervised concept discovery in generative music models. We show how sparse autoencoders applied to MusicGen residual streams reliably surface interpretable features, extending recent interpretability techniques beyond text and vision into audio/music. This is the first application of SAEs in the audio domain, to our knowledge.
@@
    Empirical insight into layer- and model-size effects. We show how concept interpretability and distinctiveness vary systematically with layer depth, sparsity/expansion settings, and model scale, contributing controlled evidence to the study of feature localization in generative models.
@@ 2.1 Statistical Learning of Musical Structure
Computational approaches to modeling music span rule-based systems, probabilistic models, and modern neural networks. Early Markov models captured local dependencies like chord transitions but failed at longer-range structure (Conklin, 2003), while information-theoretic analyses have revealed statistical bases for harmonic categories (Jacoby et al., 2015). For several years, neural architectures have achieved impressive results in both symbolic and audio modeling (Huang et al., 2018; Van Den Oord et al., 2016; Mehri et al., 2016). More recent systems often model audio directly with transformers or diffusion approaches (Agostinelli et al., 2023; Copet et al., 2024; Liu et al., 2023; Huang et al., 2023). Yet as these models generate increasingly sophisticated music, our understanding of what musical knowledge they acquire lags behind. The progress in generation has coincided with greater opacity of internal structure, motivating our central question: what do these models actually learn about music? Our work seeks to make progress on answering this question.
@@ 2.4 Automated Concept Discovery and Evaluation
A central challenge in interpretability is to discover concepts a model has learned organically, rather than imposing them a priori. Recent frameworks address this across domains. InterPLM (Simon & Zou, 2024) trains sparse autoencoders on a protein language model (ESM-2), uncovering thousands of latent features, many aligning with known biological concepts and others appearing novel. Concept Bottleneck Models (CBMs) (Koh et al., 2020) instead constrain intermediate layers to discrete, named concepts, though typically with a hand-specified set. Recent variants invert this paradigm, using SAEs to first discover latent concepts and then name them automatically or semi-automatically (Rao et al., 2024). While developed outside of music, these principles are broader: generative music models likely encode diverse concepts without explicit labels. Automated discovery and evaluation of such features could reveal how these models “understand” music, and enable finer control over generated outputs.
@@ 3.1 Dataset and Activation Extraction
The first step in our pipeline is to obtain a dataset of music clips for which we can obtain activations from the generative model. We use the MusicSet (Wei et al., 2024b) dataset for this, which is a collection of around 160,000 samples, most of which are $\approx$10 seconds long. MusicSet is a combination of data from MTG-Jamendo (Bogdanov et al., 2019), MusicCaps (Agostinelli et al., 2023), and MusicBench (Melechovský et al., 2023). All the source datasets are Creative Commons-licensed (though only MTG-Jamendo explicitly states that this applies to the audio). We chose this dataset because it offers a comparatively diverse range of genres, instrumentation, and musical characteristics at a larger scale than many alternatives. We also found the data to be more stylistically diverse and high-quality than FMA (Defferrard et al., 2016), another popular option. We produce activations by feeding the music into two pre-trained MusicGen (Copet et al., 2024) models: MusicGen-Large (henceforth, MGL) and MusicGen-Small (henceforth, MGS). For each, we extract activation vectors from five residual stream layers: the early layer (layer 2), late layer (second-to-last), middle layer, and layers at 25%, 50%, and 75% depth. This corresponds to $\textrm{{Layer}}_{\textrm{MGS}}\in\{2,6,12,18,22\}$, and $\textrm{{Layer}}_{\textrm{MGL}}\in\{2,12,24,36,46\}$. Activation dimensionality is 1024 for MGS and 2048 for MGL. [^2] Note: we use unconditional audio; see Appendix C for more details.
@@ 3.5 Automated Interpretability
    A generative label proposal strategy. We query a large multimodal model (i.e., Gemini Flash 1.5, but explored others in Appendix B) with concatenated audio of the top-10 max activating examples for each feature. We instruct the model to produce a set of concept tags, confidence scores for each, and an overarching label and description for the feature.
@@
    A classifier-based strategy where we extract a set of tags for each feature using pre-trained Essentia (Bogdanov et al., 2013; Alonso-Jiménez et al., 2020) audio classifier models. We consider multiple top predicted tags from each model as candidate labels.
@@ 3.6 Human Validation
Ultimately, given the complexity of musical signals and the possible ambiguity of features, we conducted a human validation study (IRB approved) to assess labeling quality. For each feature, participants were presented with three audio examples (sampled from SAEs with high valid feature counts) and the candidate labels. They selected the label that best captured the commonality across the examples, rated their confidence in that choice, and provided an optional free-text label.
@@ 3.7 Generation Steering
Experimental Setup. We evaluate steering using a neutral prompt, ”Simple melody” (as in Arad et al. (2025)) for steering strengths $\alpha\in\{0.0,1.0\}$. For each feature, we generate an unsteered baseline ($\alpha=0$) and a maximally steered version ($\alpha=1$) to test the “steerability” of a concept. We used MusicGen-Large SAEs with exp. factor 32, $k\in\{32,100\}$, and middle to late layers (24, 36, and 46), based on feature statistics.
@@ 4.1 Statistics of Discovered Features
Before filtering, feature activations show a heavy-tailed distribution: some fire in many tracks, while most activate only rarely (see Figure 6 for details). SAE hyperparameters appear to mediate this balance, i.e. larger $k$ or expansion factor ($EF$) increase overall recovery but also changes the activation frequency distributions. This imbalance makes raw feature counts misleading and motivates our explicit filtering approach, to isolate features that are both selective and interpretable.
@@
Table 1: Feature counts across models, layers, and SAE configurations (total 4697). For each model layer, the configuration providing the max. num. of features is bolded.
@@ 4.6 Steering with Discovered Concepts
Table 2 summarizes results. Depending on SAE configuration, between 15–35% of tested features showed improved CLAP alignment with their top-activating examples under steering, starting from the prompt “Simple Melody.” While these proportions may appear modest, they demonstrate that a non-trivial fraction of discovered features correspond to causal, manipulable directions in activation space even under the automatic labeling scheme. This shows that interpretable features are not only descriptive, but in many cases actionable for controlling outputs. Figure 5 illustrates such cases, where steering shifts outputs toward the intended feature class while holding prompt and seed fixed. This establishes the potential for SAE-driven steering in controllable generation settings, though our primary goal in this work remains feature discovery (Peng et al., 2025).
@@ Acknowledgments
We gratefully acknowledge AWS for providing computational resources through the AWS Research and Engineering Studio (RES) pilot program. We thank Brian McCarthy and Jared Novotny for their valuable support, and Peter Fisher (MIT ORCD) for making this collaboration happen. MC is supported by a fellowship from “la Caixa” Foundation (ID 100010434) with code LCF/BQ/EU23/12010079. We also thank Michael Casey, Anna Huang, and the HAI-Res lab for supportive comments.
@@ Appendix B Choice of Captioning Model
Appendix C Prompt-Conditioned Activations
@@
Appendix D Instructions Given to Participants

### v3

# Discovering and Steering Interpretable Concepts in Large Generative Music Models

@@ Abstract
The fidelity with which neural networks can now generate content such as music presents a scientific opportunity: these systems appear to have learned implicit theories of such content’s structure through statistical learning alone. This offers a potentially new lens on theories of human-generated media. When internal representations align with traditional constructs (e.g. chord progressions in music), they show how such categories can emerge from statistical regularities; when they diverge, they expose limits of existing frameworks and patterns we may have overlooked but that nonetheless carry explanatory power. In this paper, focusing on autoregressive music generators, we introduce a method for discovering interpretable concepts using sparse autoencoders (SAEs), extracting interpretable features from the residual stream of a transformer model. We make this approach scalable and evaluable using automated labeling and validation pipelines. Our results reveal both familiar musical concepts and coherent but uncodified patterns lacking clear counterparts in theory or language. As an extension, we show such concepts can be used to steer model generations. Beyond improving model transparency, our work provides an empirical tool for uncovering organizing principles that have eluded traditional methods of analysis and synthesis.
Figure 1: Multi-stage pipeline for discovering and steering interpretable concepts in autoregressive music models. (1) Music from a large corpus is passed through a pre-trained generator to extract activations from multiple layers. (2) Sparse autoencoders reconstruct activations (also usable for steering), and features are filtered to retain the most viable candidates. (3) Retained features are characterized via musical examples and labeled using generative labeling with a multimodal LM and classifier-based labeling with pre-trained models.
@@ 1 Introduction
    A general pipeline for unsupervised concept discovery in autoregressive music models. We show how sparse autoencoders applied to MusicGen residual streams reliably surface interpretable features, extending recent interpretability techniques beyond text and vision into audio/music. This is the first application of SAEs in audio, to our knowledge.
@@
    Empirical insight into layer- and model-size effects. We show how concept interpretability and distinctiveness vary with layer depth, sparsity/expansion settings, and model scale, contributing evidence to the study of feature localization in generative models.
@@ 2.1 Statistical Learning of Musical Structure
Computational approaches to modeling music span rule-based systems, probabilistic models, and modern neural networks. Early Markov models captured local dependencies like chord transitions but failed at longer-range structure (Conklin, 2003), while information-theoretic analyses have revealed statistical bases for harmonic categories (Jacoby et al., 2015). For several years, neural architectures have achieved impressive results in both symbolic and audio modeling (Huang et al., 2018; Van Den Oord et al., 2016; Mehri et al., 2016). More recent systems often model audio directly with transformers or diffusion approaches (Agostinelli et al., 2023; Copet et al., 2024; Liu et al., 2023; Huang et al., 2023). Yet as these models generate increasingly sophisticated music, our understanding of what musical knowledge they acquire lags behind. The progress in generation has coincided with greater opacity of internal structure, motivating our central question: what do these models actually learn about music? Our work seeks to make progress on answering this question (see Appendix G for additional music-theoretic context).
@@ 2.4 Automated Concept Discovery and Evaluation
A central challenge in interpretability is to discover concepts a model has learned organically, rather than imposing them a priori. Recent frameworks address this across domains. Both Simon & Zou (2024) and Gujral et al. (2025) train sparse autoencoders on a protein language model (ESM-2), uncovering thousands of latent features, many aligning with known biological concepts and others appearing novel. Concept Bottleneck Models (CBMs) (Koh et al., 2020) instead constrain intermediate layers to discrete, named concepts, though typically with a hand-specified set. Recent variants invert this paradigm, using SAEs to first discover latent concepts and then name them automatically or semi-automatically (Rao et al., 2024). While developed outside of music, these principles are broader: generative music models likely encode diverse concepts without explicit labels. Automated discovery and evaluation of such features could reveal how these models “understand” music, and enable finer control over generated outputs.
@@ 3.1 Dataset and Activation Extraction
The first step in our pipeline is to obtain a dataset of music clips for which we can obtain activations from the generative model. We use the MusicSet (Wei et al., 2024b) dataset for this, which is a collection of around 160,000 samples, most of which are $\approx$10 seconds long. MusicSet is a combination of data from MTG-Jamendo (Bogdanov et al., 2019), MusicCaps (Agostinelli et al., 2023), and MusicBench (Melechovský et al., 2023). All the source datasets are Creative Commons-licensed (though only MTG-Jamendo explicitly states that this applies to the audio). We chose this dataset because it offers a comparatively diverse range of genres, instrumentation, and musical characteristics at a larger scale than many alternatives. We also found the data to be more stylistically diverse and high-quality than FMA (Defferrard et al., 2016), another popular option. We produce activations by feeding the music into two pre-trained MusicGen (Copet et al., 2024) models: MusicGen-Large (henceforth, MGL) and MusicGen-Small (henceforth, MGS). For each, we extract activation vectors from five residual stream layers: the early layer (layer 2), late layer (second-to-last), middle layer, and layers at 25%, 50%, and 75% depth. This corresponds to $\textrm{{Layer}}_{\textrm{MGS}}\in\{2,6,12,18,22\}$, and $\textrm{{Layer}}_{\textrm{MGL}}\in\{2,12,24,36,46\}$. Activation dimensionality is 1024 for MGS and 2048 for MGL. Note: we use unconditional audio; see Appendix E for more details.
@@ 3.5 Automated Interpretability
    A generative label proposal strategy. We query a large multimodal model (i.e., Gemini Flash 1.5, but explored others in Appendix B) with concatenated audio of the top-10 max activating examples for each feature (see prompt in Appendix H). We instruct the model to produce a set of concept tags, confidence scores for each, and an overarching label and description for the feature.
@@
    A classifier-based strategy where we extract a set of tags (see Appendix D) for each feature using pre-trained Essentia (Bogdanov et al., 2013; Alonso-Jiménez et al., 2020) audio classifier models. Multiple top predicted tags from each model are candidate labels.
@@ 3.6 Human Validation
Ultimately, given the complexity of musical signals and the possible ambiguity of features, we conducted a human validation study (IRB approved) to assess labeling quality. For each feature, participants were presented (see Appendix F) with three audio examples (sampled from SAEs with high valid feature counts) and the candidate labels. They selected the label that best captured the commonality across the examples, rated their confidence in that choice, and provided an optional free-text label.
@@ 3.7 Generation Steering
Experimental Setup. We evaluate steering using a neutral prompt ”Simple melody” (as in Arad et al. (2025)) for steering strengths $\alpha=0.0$ (unsteered baseline) and $\alpha=1.0$ (max. steered version) for each feature to test the “steerability” of a concept. We used MusicGen-Large SAEs with exp. factor 32, $k\in\{32,100\}$, and middle to late layers (24, 36, and 46), based on feature statistics.
@@ 4.1 Statistics of Discovered Features
Initially, feature activations show a heavy-tailed distribution: some fire in many tracks, while most activate only rarely (see Figure 6 for details). SAE hyperparameters appear to mediate this balance, i.e. larger $k$ or expansion factor ($EF$) increase overall recovery but also changes the activation frequency distributions. This imbalance makes raw feature counts misleading and motivates our explicit filtering approach, to isolate features that are both selective and interpretable.
@@
Table 1: Feature counts across models, layers, and SAE configurations (total 4697). For each model layer, the configuration providing the max. number of features is bolded.
@@ 4.4 Concept Representation
How do features co-activate?
In Appendix C, we analyze feature co-activation patterns, showing examples of cross-layer concept connections which emerge from SAE training.
@@ 4.6 Steering with Discovered Concepts
Table 2 summarizes results. Depending on SAE configuration, 15–35% of tested features had improved CLAP alignment with their top-activating examples under steering (with prompt “Simple Melody”). While this may appear modest, it demonstrates that a non-trivial fraction of discovered features correspond to causal, manipulable directions in activation space even under automatic labeling: actionable for controlling outputs. Figure 5 illustrates such cases, where steering shifts outputs toward the intended feature class while holding prompt and seed fixed. This establishes the potential for SAE-driven steering in controllable generation settings, though our primary goal in this work remains feature discovery (Peng et al., 2025).
@@
We conducted a listening study with 10 participants to each listen to 10 sets of clips; each containing a baseline, a steered version, a random-direction matched-norm steered version, and a representative of the steering target, asking them to match the candidates (baseline, randomly steered, and SAE steered). Each participant’s 10 sets were equally sampled from the total of top 50 steerable features. Participants largely selected the SAE-steered audio (66/100, compared to 17 each for baseline and random; $\chi^{2}=48.02,\ p<.0001$). This suggests that the steering effects are clearly perceptible.
@@ Acknowledgments
We gratefully acknowledge AWS for providing computational resources through the AWS Research and Engineering Studio (RES) pilot program. We thank Brian McCarthy and Jared Novotny for their valuable support, and Peter Fisher (MIT ORCD) for making this collaboration happen. The authors acknowledge the MIT Office of Research Computing and Data for providing high performance computing resources that have contributed to the research results reported within this paper. MC is supported by a fellowship from “la Caixa” Foundation (ID 100010434) with code LCF/BQ/EU23/12010079. We also thank Michael Casey, Anna Huang, and the HAI-Res lab for supportive comments.
@@ Appendix B Choice of Captioning Model
Appendix C Feature Co-activations
We conducted large-scale measurements of feature co-activation across models, layers, and SAE configurations in order to understand when distinct learned features might activate on the same musical input. For a pair of features $(f_{i},f_{j})$, we treat the number of tracks common among their respective top-10 activating examples as an indicator of functional relatedness. Although the most common non-zero outcome is that two features overlap on exactly one track, the long tail of rarer, higher-overlap cases provides (observational) evidence of meaningful structure. Here we summarize the most clear and recurring qualitative patterns observed in this tail.
Same-site convergence across distinct SAEs.
First, we find cases where separate SAEs trained on the same model recover nearly identical concepts at the same layer position. Representative examples include:
  • Distorted Rock (MGL L36) by three separate SAE configurations: $k=32,EF=32;k=100,EF=32;k=100,EF=4$
  • String Orchestra (MGL L2) by $k=100,EF=4;k=100,EF=32$
  • Reverberant Vocals and Percussion (MGS L22) by $k=100,EF=32;k=32,EF=32$
These patterns indicate that certain musical concepts may be encoded robustly enough in the model’s residual stream that different SAEs can reliably recover them.
Cross-layer hierarchical relationships.
We also observe cases where early-layer features and deeper-layer features appear to express related musical ideas at different (e.g. increasing) levels of abstraction. For example:
  • Wind-Dominated Folk Drone (L12) $\to$ Eastern European Folk Wind Music (L36) in MGL $k=32,EF=32$ with 50% overlap between features (5/10 examples)
  • Jazzy Brass (L24) $\to$ Brass and Syncopation (L45) in MGL $k=100,EF=4$ with 50% overlap
  • Arpeggiated Electronic Music (L6) $\to$ Synth Pop (L18) in MGS $k=32,EF=32$ with 40% overlap
Such patterns point to potentially semantically hierarchical pathways in the learned features, where deeper layers might develop earlier features.
Overall structure.
Across the full set of experiments, all configurations of MGL and MGS exhibit the same broad pattern: most feature pairs co-activate sparsely if at all (see right side of Figure 10). Figure 11 shows statistics of features co-activating across layers. These results suggest that certain SAE configurations are more successful at recovering these than others. For the most successful one, we show layer co-activation statistics in the heatmap to the right (within-layer co-activations are zeroed out here). Co-activations are distributed rather than e.g. locally biased. Finally, in Figure 12, we examine to what degree different SAE configurations can recover similar features in similar locations. Here, there is a split between cooccurrence rates and strengths: MGL L2 has by far the most cooccurring features across different SAE configurations, but the lowest average strength. In contrast, MGS L22 has among the fewest cooccurring features, but the highest average strength. This could be due to, for instance, the degree of compression within the model and layer: L22 is the latest layer we test in the small-scale MGS and may have more well-formed features that are consistently recovered, while L2 is the earliest layer we test in the large-scale MGL, and could have more generic and low-level features. We note however that these are observations rather than formal, mechanistic findings.
Figure 10: Co-activation/cooccurrence statistics by unit of analysis: between SAE configurations on the same model/layer, between layers on the same model and SAE configuration, within the same layer/model/SAE configuration, and between MGS and MGL models.
Figure 11: Statistics of cross-layer co-activations.
Figure 12: Statistics of co-activations across SAE configurations, within the same model and layer.
In aggregate, these observations suggest that while the bulk of discovered features behave independently, there is a nontrivial amount of structure learned across different parts of the models that co-activation analysis can help to reveal. A deeper analysis might look at how these observations change under different co-activation strength thresholds, however we acknowledge this requires a much larger scope of experimentation.
Appendix D Essentia Tags
Essentia Model Tags
We use the following Essentia feature extractors and tagging models. By default, we select the top 3 predicted tags from each model, since we observe too much noise in the results to rely solely on the top predicted tag and these are often used in multi-label settings (exceptions are noted and justified below):
  • Genre-Discogs400: Genre prediction (top 3)
  • Mood-Jamendo: Multi-label mood/theme (top 3)
  • Instrument-Jamendo: Instrument presence (top 4, to account for common band compositions)
  • Tags-Jamendo: General music tags (top 3)
  • Tags-MSD: General music tags, different set (top 3)
  • Moods-Audioset: Mood categories, different set (top 2; these tend to be less distinct)
These in turn depend on VGGish, EffNet, or MusicCNN embeddings, all of which are also extracted.
Appendix E Prompt-Conditioned Activations
@@
Appendix F Instructions Given to Participants